OpenStudy (anonymous):

Help please Use logarithmic differentiation to find the derivative of y=(cosx)^tanx

OpenStudy (tanya123):

123, ABC, 098, ZYX

OpenStudy (tanya123):

Call me ;) ^

OpenStudy (fibonaccichick666):

so, if I understand the terminology, this wants you to use properties of logs to make differentiation easier

OpenStudy (fibonaccichick666):

so, what do you think your first step should be? \[y=(cosx)^{tanx}\]

OpenStudy (fibonaccichick666):

@sasha.o

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle y=f(x)^{g(x)} }\) To find the derivative (dy/dx), this is what you do: (Abstract example, for some differentiable functions f and g) \(\bf \color{red}{[1]}\) Hit both sides with a natural log, ln. \(\large\color{black}{ \displaystyle \ln\left(y\right)=\ln\left(f(x)^{g(x)}\right) }\) \(\bf \color{red}{[2]}\) Use \(\ln(A^B)=B\ln(A)\). \(\large\color{black}{ \displaystyle \ln\left(y\right)={g(x)}\ln\left(f(x)\right) }\) \(\bf \color{red}{[3]}\) Differentiate both sides, and remember that y will get a chain rule of y' since y is really here to denote a function of x. And for the left side you get use product rule \(dy/dx\)(f•g)=f\('\)g+f g\('\) \(\large\color{black}{ \displaystyle \frac{y'}{y}=g'(x)\ln\left(f(x)\right)+g(x)\frac{f'(x)}{f(x)} }\) \(\large\color{black}{ \displaystyle \frac{y'}{y}=g'(x)\ln\left(f(x)\right)+\frac{f'(x)g(x)}{f(x)} }\) \(\bf \color{red}{[4]}\) Multiply both sides times y. \(\large\color{black}{ \displaystyle y'=y\left[g'(x)\ln\left(f(x)\right)+\frac{f'(x)g(x)}{f(x)}\right] }\) \(\bf \color{red}{[5]}\) You are given the function y in the beginning. \(\large\color{black}{ \displaystyle y'=f(x)^{g(x)}\left[g'(x)\ln\left(f(x)\right)+\frac{f'(x)g(x)}{f(x)}\right] }\)